Circular No. 8346

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 H6 (SWAN) On May 13, the Central Bureau was informed apparently indepen-

dently by X.-m. Zhou, K. Černis, and M. Mattiazzo of the existence of a possible unknown comet on the low-resolution ultraviolet SWAN images taken from the SOHO spacecraft. As there have been numerous neverconfirmed reports to the Central Bureau of such objects over the last couple of years, despite a great exertion of effort to have them confirmed by ground-based observers, we proceeded with caution to try and get solid confirming evidence of a real comet. Reports trickled in (as can be seen from the table below), but successive attempts to produce search ephemerides met with some negative results that indicated some of the early positions were apparently not of the new comet.

2004 UT  α_{2000}  δ_{2000}  Mag. Observer
Apr. 29  4^{h}09^{m}  − 9^{o}.4  SWAN
May 2  4 05  −10.3  _{”}
5  3 59  −10.6  _{”}
8  3 56  −11.4  _{”}
14.388  3 45 56^{s}  −11 47′.5  7.8 Mattiazzo
14.444  3 46.0  −11 44  7.9 Pearce
16.37777 3 42 51.52 − 12 01 01′′.7  9.2 Mattiazzo
21.84565 3 33 47.86 − 12 37 44.4  _{”}
21.85200 3 33 47.24 − 12 37 46.1  _{”}
21.85689 3 33 46.57 − 12 37 47.7  _{”}
21.86298 3 33 46.46 − 12 37 50.1  _{”}
22.807 3 32 18 − 12 41.2 7.5 Drescher
25.84159 3 26 56.59 − 12 57 13.4  Mattiazzo
25.84900 3 26 55.70 − 12 57 13.5  _{”}
25.85747 3 26 54.98 − 12 57 15.4  _{”}
25.86269 3 26 54.43 − 12 57 16.5  _{”}

M. Mattiazzo (Wallaroo, S. Australia). Visually with 25×100 binoculars on May 14; coma diameter 3′. CCD astrometry: May 16, 135-mm telephoto lens. May 21, λ = 138^{o}40′58′′, ϕ = −31^{o}08′26′′; narrow tail > 8′ long in p.a. 187^{o} (0.28-m reflector). May 25, λ = 135^{o}.550, ϕ = −34^{o}.926 (0.28-m refl.).

A. Pearce (Nedlands, W.A.). Visual, 0.20-m refl. Moderately condensed. C. Drescher (Leyburn, Qld.). 40×100 binoculars; diffuse coma of diameter

1′.5 with some condensation.

2004 May 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green